40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6634     Entry time: Wed May 9 14:32:31 2012
Author: Jenne 
Type: Update 
Category: CDS 
Subject: Burt restored 

Den and Alex left things not-burt restored, and Den mentioned to me that it might need doing.

I burt restored all of our epics.snaps to the 1am today snapshot.  We lost a few hours of striptool trends on the projector, but now they're back (things like the BLRMS don't work if the filters aren't engaged on the PEM model, so it makes sense).

ELOG V3.1.3-